OpenStudy (anonymous):

plz , need some help :) will fan+medal Sketch the following sets and determine which are domains: (a) |z − 2 + i| ≤ 1; b) |2z + 3| > 4; (c) Imz > 1; (d) Im z = 1; (e) 0 ≤ arg z ≤ π/4 (z = 0); (f) |z − 4| ≥ |z|.

OpenStudy (anonymous):

YOu are working complex plane right?

OpenStudy (anonymous):

yeah, but dnt knw anything

OpenStudy (anonymous):

and besides your questino is incomplete.. if its domain/..??

OpenStudy (anonymous):

hmm ok ill post the whol qs

OpenStudy (anonymous):

Sketch the following sets and determine which are domains: (a) |z − 2 + i| ≤ 1; b) |2z + 3| > 4; (c) Imz > 1; (d) Im z = 1; (e) 0 ≤ arg z ≤ π/4 (z = 0); (f) |z − 4| ≥ |z|.

OpenStudy (anonymous):

i dnt knw how to graph or what does domain mean in complex plane i missed te class XD

terenzreignz (terenzreignz):

Someone's not a very good student :P

OpenStudy (anonymous):

yeah thats me lol XD

terenzreignz (terenzreignz):

I was referring to me, but okay :D

OpenStudy (anonymous):

i got 2 warning for not going to my classes hehe

OpenStudy (anonymous):

@terenzreignz lol ok XD

OpenStudy (anonymous):

besides.. don't you substitude z = x + iy ?

OpenStudy (anonymous):

i dnt think so its the same to complex plane XD

terenzreignz (terenzreignz):

No offense, Mashy, but that's not the meaning of domain in this context. http://en.wikipedia.org/wiki/Domain_(mathematical_analysis)

OpenStudy (anonymous):

see this is ive got Ans. (b), (c) are domains.

OpenStudy (anonymous):

but dnt knw how :o

terenzreignz (terenzreignz):

Magic. How did you get those answers, anyway? :P

OpenStudy (anonymous):

ermm.. dang.. my complex math is rusty then :-/

OpenStudy (anonymous):

there is some magic book called "manual" for cheating :P

OpenStudy (anonymous):

thanks for pointing it out.. ll delete that embarrassing post immediately :-/

OpenStudy (anonymous):

no lol ugh

OpenStudy (anonymous):

yea no he is right.. if someone read this thread.. that post would only confuse them!!!

OpenStudy (anonymous):

be the man , and let ur words :P

OpenStudy (anonymous):

terrenz.. be the man.. and help him solve :P

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

help her*

OpenStudy (anonymous):

anyway gtg nw XD i hope to get it somtime ltr XD

OpenStudy (anonymous):

ty for help @terenzreignz @Mashy

OpenStudy (anonymous):

sorry.. for that gender mistake :P i ll try first one (a) |z − 2 + i| ≤ 1; \[z = x +iy\] \[|(x-2)+i(y+1)|\le 1\]\[(x-2)^{2} + (y+1)^2 \le 1^2 or 1\] so now can you sketch that? on a complex plane? do you recognize that oddly familiar equation?

OpenStudy (anonymous):

circl ?

OpenStudy (anonymous):

bingo.. u see this is just like ordinary math, just that y axis is the imaginary axis.. (it goes like 0, i, 2i, 3i, .. ) multiples of i :D

OpenStudy (anonymous):

oh , ic ok ill try the rest then post them gtg nw bye ^^

OpenStudy (anonymous):

oki doki.. see ya!

terenzreignz (terenzreignz):

Also, save yourself the trouble... Granted that Mashy has pretty much demonstrated it, regardless... Let me give you a hint... and a very basic and fundamental one, at that. \[\Large |z - z_o| = k\] ALWAYS MEANS A circle centred around \(\large z_o\) with radius k.

OpenStudy (anonymous):

it means uper circle ^^